OpenStudy (anonymous):

can someone help me factor 4x^2+20x-24?

OpenStudy (anonymous):

\[ 4x^2 +20x -24 = 4 \cdot (x^2 + 5x - 6) = 4 \cdot (x-1)(x+6) \]